OpenStudy (mathmath333):

There are five boxes in a cargo hold. The wt of the first box is 200 kg, 2nd is 20%higher than the wt of the 3rd whose wt is 25% higher than the wt of the first box. The 4th box is 350 kgs and is 30% lighter than the fifth box. find the difference In the avg wt of the heaviest 3 and the lightest 3?

OpenStudy (mathmath333):

options 1.).116.66 2.) 125 c.) 150 d.) 112.5

OpenStudy (mathmath333):

weight of the boxes i found 1st box=200 kg 2nd box=300 kg 3rd box=250 kg 4th box=350 kg 5th box=500 kg

OpenStudy (mathmath333):

@gorv

OpenStudy (loser66):

Hand off, no matter what the way I go, I get 133.33 which is non of the options.

OpenStudy (mathmath333):

i too got 133.3 but book says 116.66 is correct

OpenStudy (mathmath333):

@amistre64

OpenStudy (amistre64):

There are five boxes in a cargo hold. The wt of the first box is 200 kg, 2nd is 20%higher than the wt of the 3rd: 200(1.25)(1.2) 3rd whose wt is 25% higher than the wt of the first box: 200(1.25) The 4th box is 350 kgs and is 30% lighter than the fifth box. 350 = f(.70) find the difference In the avg wt of the heaviest 3 and the lightest 3 200 300 250 350 500 (200+300+250)/3 - (350+500)/2

OpenStudy (amistre64):

pfft, 3 and 3 ...

OpenStudy (amistre64):

133.3 so we might be interpreting the question incorrectly, or they simple gave a wrong solution

OpenStudy (mathmath333):

this one book gave hint (350+500)-(300+200)=350 and 350/3=116.66

OpenStudy (amistre64):

isnt the 300 the common between them?

OpenStudy (amistre64):

(a+b+c) - (x+y+c) = a+b-x-y

OpenStudy (texaschic101):

heaviest 3 .... (500 + 350 + 300) / 3 = 1150/3 = 383.33 lightest 3 ... (200 + 250 + 300) / 3 = 750/3 = 250 383.33 - 250 = 133.33 ...oops that didn't help

OpenStudy (amistre64):

i saw a solution manual say 0-2 = 4 so im not impressed with book solutions

OpenStudy (mathmath333):

so is 133 the final vedict

OpenStudy (amistre64):

as long as we have the wghts correct :0 then yes

OpenStudy (mathmath333):

ok thanks every one i was worried cuz this book is of 5th edition so i thought i may be wrong